Solve using the elimination method 10a+6b=8
5a+3b=2
Thank you for your assistance

Answer by ewatrrr(24785)

 
Hi,
Solve using the elimination method
10a+6b=8
5a+3b=2
10a+6b = 8
-10a-6b =-4 |multiplying 2nd Eq thru by -2
0 ≠ 4 there is NO SOLUTION